OpenStudy (anonymous):

Here are the vertices of my regular tetrahedron, taken from a cube centered at (0,0,0) \[A:=(1,1,1)\] \[B:=(−1,1,−1)\] \[C:=(−1,−1,1)\] \[D:=(1,−1,−1)\] what is the area of one face of that tetrahedron?

OpenStudy (anonymous):

The area corresponds to the determinant, or to the cross product, between 2 adjacent edge vectors.

OpenStudy (anonymous):

I will consider 2 adjacent vectors\[AB = (-2,0,-2)\] \[AC = (-2,-2,0)\]

OpenStudy (anonymous):

what is their cross product?

OpenStudy (anonymous):

it is det( i j k -2 0 -2 -2 -2 0 ) which is -4(i-j-k)

OpenStudy (anonymous):

or (-4, 4, 4) what is the length of that vector

OpenStudy (anonymous):

\[4\sqrt{3}\] what is half of that? \[2\sqrt{3}\] that is the area of one face of my regular tetrahedron taken from a cube with side length 2

OpenStudy (anonymous):

is that right? I think I am wrong

OpenStudy (anonymous):

just find the magnitude of say AB then use trig or pythagorean theorem to find the height of a face then do 1/2 base times height

OpenStudy (anonymous):

the length of (-4,4,4) would be sqrt(4^2+4^2+4^2)

OpenStudy (anonymous):

hmm... \[12 = 24 - 4\sqrt{3}*\frac{1}{2}\]doesn't sound right

OpenStudy (anonymous):

but that's because I wrote it wrong. \[12=24-2(4*3)*0.5\]

OpenStudy (anonymous):

yeah, \[2\sqrt{3}\]is the area of the side! thanks

OpenStudy (anonymous):

It's a regular tetrahedron, so the faces are all equilateral triangles. The sides of the triangles are diagonals of the cube faces, so sqrt(8). In an 30-60-90 right triangle (half of an equilateral triangle) the sides are in the ratio 1-sqrt(3)-2. So the height of one of the triangular faces will be sqrt(3)*sqrt(8)/2 = sqrt(6). Then the area will be 1/2sqrt(6)sqrt(8) = 1/2sqrt(48) = 2sqrt(3). The total area of the tetrahedron must be less than the total are of the cube. 4(2sqrt(3)) = 8*1.732 = 13.856 6(4) = 24 So that checks out.